# Taming singularities and chaos in conformal gravity

### v1

@@ I Introduction
As a follow up of several previous works [1, 2, 3, 4, 5, 6, 7, 8, 9], we hereby address the singularity issue in theories enjoying Weyl conformal invariance, whose crucial role was pointed out for the first time in Ref. [10] and more recently in Ref. [11]. In the past, we showed how the Weyl invariance can be implemented in order to make spherically symmetric or axisymmetric spacetimes geodetically complete [1]. We also solved the big bang singularity problem [12, 13] and, more recently, we discovered a class of stable wormhole solutions sourced by normal not exotic matter [14]. Moreover, a proper and unique conformal frame can also explain the galactic rotation curves [15].
@@ II Bianchi IX Spacetime
$\omega^{1}=dx+\sin ydz\,$, (3)
$\omega^{2}=\cos xdy-\sin x\cos ydz\,$,
$\omega^{3}=\sin xdy+\cos x\cos ydz\,$,
@@
$\xi_{1}=\left(0,\ \sec y\cos z,\ \sin z,\ -\tan y\cos z\right),$ (6)
$\xi_{2}=\left(0,\ -\sec y\sin z,\ \cos z,\ \tan y\sin z\right),$
$\xi_{3}=\left(0,\ 0,\ 0,\ 1\right).$
@@
$\begin{aligned} &a_{1}(t)=\exp\left(\alpha(t)+\beta_{+}(t)+\sqrt{3}\beta_{-}(t)\right),\\ &a_{2}(t)=\exp\left(\alpha(t)+\beta_{+}(t)-\sqrt{3}\beta_{-}(t)\right),\\ &a_{3}(t)=\exp\left(\alpha(t)-2\beta_{+}(t)\right).\end{aligned}.$ (9)
@@ II.1 Dynamics of the Bianchi IX metric
$\left(\frac{d\beta_{+}}{d\eta}\right)^{2}+\left(\frac{d\beta_{-}}{d\eta}\right)^{2}$ $\leq-\frac{1}{2}V+\left(\frac{d\alpha}{d\eta}\right)^{2}_{max}-\frac{8\pi}{3}\left(e^{2\alpha}\rho\right)_{max}$ (21)
$\leq\frac{1}{4}+\left(\frac{d\alpha}{d\eta}\right)^{2}_{max}-\frac{8\pi}{3}\left(e^{2\alpha}\rho\right)_{max}\,.$
@@ III.2 Geodesic completion: massless particles
$f_{1}(x,y,z;E_{i})=\cos y\left(E_{1}\cos z-E_{2}\sin z\right)+E_{3}\sin y\,$, (41)
$f_{2}(x,y,z;E_{i})=\cos x\left(E_{1}\sin z+E_{2}\cos z\right)-\sin x\left[\sin y\left(E_{2}\sin z-E_{1}\cos z\right)+E_{3}\cos y\right],$
$f_{3}(x,y,z;E_{i})=\cos x\left[\sin y\left(E_{2}\sin z-E_{1}\cos z\right)+E_{3}\cos y\right]+\sin x\left(E_{1}\sin z+E_{2}\cos z\right).$
@@ III.3 Geodesic completion: conformally coupled and massive particles
$\dot{x}$ $=\cos y\frac{F_{1}\cos z-F_{2}\sin z+F_{3}\tan y}{a_{1}^{2}(t)\left(C\phi+m\right)}$ (56)
$+\cos y\frac{\sin x\tan y(\cos x\sec y(F_{1}\sin z+F_{2}\cos z)-\sin x(-F_{1}\tan y\cos z+F_{2}\tan y\sin z+F_{3}))}{a_{2}^{2}(t)\left(C\phi+m\right)}$
$-\cos y\frac{\cos x\tan y(\cos x(-F_{1}\tan y\cos z+F_{2}\tan y\sin z+F_{3})+\sin x\sec y(F_{1}\sin z+F_{2}\cos z))}{a_{3}^{2}(t)\left(C\phi+m\right)}$
$\dot{y}$ $=\frac{-2\sin x\cos x\sin y\left(a^{2}_{2}(t)-a^{2}_{3}(t)\right)(F_{1}\cos z-F_{2}\sin z)+F_{3}\sin(2x)\cos y\left(a_{2}(t)^{2}-a_{3}(t)^{2}\right)}{a_{2}(t)^{2}a_{3}(t)^{2}\left(C\phi+m\right)}$ (57)
$+\frac{2(F_{1}\sin z+F_{2}\cos z)\left(a_{2}(t)^{2}\sin^{2}x+a_{3}(t)^{2}\cos^{2}x\right)}{a^{2}_{2}(t)a^{2}_{3}(t)\left(C\phi+m\right)}$
$\dot{z}$ $=\frac{\cos x(\cos x(-F_{1}\tan y\cos z+F_{2}\tan y\sin z+F_{3})+\sin x\sec y(F_{1}\sin z+F_{2}\cos z))}{a^{2}_{3}(t)\left(C\phi+m\right)}$ (58)
$+\frac{\sin x(\sin x(-F_{1}\tan y\cos z+F_{2}\tan y\sin z+F_{3})-\cos x\sec y(F_{1}\sin z+F_{2}\cos z))}{a^{2}_{2}(t)\left(C\phi+m\right)}$
@@
$\Delta\lambda_{\rm cp}$ $=\int^{T}_{0}dt\ \left[1+\frac{G_{1}^{2}}{a_{1}^{2}(t)}+\frac{G_{2}^{2}}{a_{2}^{2}(t)}+\frac{G_{3}^{2}}{a_{3}^{2}(t)}\right]^{-1/2}$ (61)
$=\int^{T}_{0}\ dt\ \frac{1}{(m+C\phi)^{2}}\left[1+\frac{\check{f}_{1}^{2}}{a_{1}^{2}(t)}+\frac{\check{f}_{2}^{2}}{a_{2}^{2}(t)}+\frac{\check{f}_{3}^{2}}{a_{3}^{2}(t)}\right]^{-1/2}\,.$
@@
$\Delta\lambda_{\rm cp}$ $=\left(m+\frac{C}{\sqrt{32\pi G}}\right)^{-2}\int_{0}^{T}dt\ \left[1+\frac{\check{f}_{1}^{2}}{a_{1}^{2}(t)}+\frac{\check{f}_{2}^{2}}{a_{2}^{2}(t)}+\frac{\check{f}_{3}^{2}}{a_{3}^{2}(t)}\right]^{-1/2}$ (62)
$\simeq\left(m+\frac{C}{\sqrt{32\pi G}}\right)^{-2}\int_{0}^{T}dt\ \left[\frac{\check{f}^{2}_{2}a_{1}^{2}(t)a_{3}^{2}(t)}{\left(a_{1}(t)a_{2}(t)a_{3}(t)\right)^{2}}\right]^{-\frac{1}{2}}$
$=\frac{1}{|\check{f}_{2}|}\left(m+\frac{C}{\sqrt{32\pi G}}\right)^{-2}\int_{0}^{T}t^{p_{2}}dt\propto T^{p_{2}+1}$
@@
$\Delta\lambda_{\rm cp}$ $=\int^{T}_{0}dt\left[\frac{1}{S(t)}+\left(\frac{{G_{1}^{*}}^{2}}{a_{1}^{2}(t)}+\frac{{G_{2}^{*}}^{2}}{a_{2}^{2}(t)}+\frac{{G_{3}^{*}}^{2}}{a_{3}^{2}(t)}\right)\right]^{-1/2}$ (65)
$=\int_{0}^{T}dt\left[\frac{1}{S(t)}+\frac{1}{\left(m+\frac{C}{\sqrt{32\pi GS(t)}}\right)^{2}}\frac{1}{S^{2}(t)}\left(\frac{{\check{f}_{1}}^{2}}{a_{1}^{2}(t)}+\frac{{\check{f}_{2}}^{2}}{a_{2}^{2}(t)}+\frac{{\check{f}_{3}}^{2}}{a_{3}^{2}(t)}\right)\right]^{-1/2}\,.$
@@
$\Delta\lambda_{\rm cp}$ $\simeq\int_{0}^{T}dt\left[\left(1+\frac{L_{1}^{2n}}{t^{2n}}\right)^{-1}+\left(1+\frac{L_{1}^{2n}}{t^{2n}}\right)^{-2}\sum_{i=1}^{3}\check{f}_{i}^{2}a_{i}^{-2}(t)\right]^{-1/2}$ (66)
$\simeq\int_{0}^{T}dt\left[\left(1+\frac{L_{1}^{2n}}{t^{2n}}\right)^{-1}+\left(1+\frac{L_{1}^{2n}}{t^{2n}}\right)^{-2}\sum_{i=1}^{3}\check{f}_{i}^{2}t^{-2p_{i}}\right]^{-1/2}$
$\simeq\int_{0}^{T}dt\left[\left(1+\frac{L_{1}^{2n}}{t^{2n}}\right)^{-1}+\left(1+\frac{L_{1}^{2n}}{t^{2n}}\right)^{-2}\check{f}_{2}^{2}t^{-2p_{2}}\right]^{-1/2}$
@@ IV Taming chaos
$\Delta p_{ij}=\!\!\!\max_{\begin{subarray}{c}0<u<1\\ i,j=1,2,3\end{subarray}}\left|p_{i}(u)-p_{j}(u)\right|=\frac{2\sqrt{3}}{3}\,.$ (70)

### v2

@@ I Introduction
As a follow up of several previous works [1, 2, 3, 4, 5, 6, 7, 8, 9], we hereby address the singularity issue in theories enjoying Weyl conformal invariance, whose crucial role was pointed out for the first time in Ref. [10] and more recently in Ref. [11]. In the past, we showed how the Weyl invariance can be implemented in order to make spherically symmetric or axisymmetric spacetimes geodetically complete [1]. We also solved the big bang singularity problem [12, 13] and, more recently, we discovered a class of stable wormhole solutions sourced by normal not exotic matter [14]. Moreover, a proper and unique conformal frame can also explain the galactic rotation curves [15] [^1] Hystorically, conformal gravity refers to the theory proposed by H. Weyl in $1918$ in order to unify gravity and the Maxwell’s electromagnetic theory [16, 17]. Such a proposal did not work, but was later resumed as a purely gravitational theory by P.D. Mannheim [18] and extended by R.K. Nesbet [19]. Such a theory is consistent with unitarity, renormalizable, and able to address issues like the galactic rotation curves and the accelerating Hubble expansion. All the results presented in this paper are based on the symmetry and on the Bianchi IX metric as an exact solution, thus they can be exported to any conformal theory sharing the same solution..
@@ II Bianchi IX Spacetime
$aligned&\omega^{1}=dx+\sin ydz\,,\\ &\omega^{2}=\cos xdy-\sin x\cos ydz\,,\\ &\omega^{3}=\sin xdy+\cos x\cos ydz\,$, (3)
@@
$aligned&\xi_{1}=\left(0,\ \sec y\cos z,\ \sin z,\ -\tan y\cos z\right),\\ &\xi_{2}=\left(0,\ -\sec y\sin z,\ \cos z,\ \tan y\sin z\right),\\ &\xi_{3}=\left(0,\ 0,\ 0,\ 1\right).$ (6)
@@
[…]
@@ II.1 Dynamics of the Bianchi IX metric
[…]
@@ III.1 Einstein’s conformal gravity
Let us here expand on the general singular nature of the rescaling. In order to remove a coordinate singularity in a diffeomorphism invariant theory, we need a singular general coordinate transformation. The reader can check about the Gullstrand-Painleve’ coordinates’ transformation, or similar. In exactly the same vain, in a Weyl conformal invariant theory we select a conformal rescaling that is singular in the regions (a point for the Schwarzschild metric, a point or two disconnected points for the FRW spacetime, a ring for the Kerr spacetime, etc.) where the geodesic equation ceases to be valid. Equivalently, geodesics cannot be extended in such regions and the predictability is lost. This procedure mutates an incomplete spacetime into a geodetically complete one making the singularities unreachable. If we choose a conformal rescaling singular in other points or regions, where the metric is perfectly fine, other physical properties can emerge, like for example the splitting of the spacetime into multiple geodetically complete regions [5, 9, 38]. In this respect the causal structure can be drastically affected by a Weyl transformation: casually connected regions can be made causally disconnected by a singular conformal transformation. However, in this paper the geodesic completion of the spacetime is achieved by means of a rescaling that is singular at the point $t=0$ that cannot be reached in finite proper time or finite affine time. Therefore, the causal structure is preserved everywhere besides $t=0$ that becomes the past infinity.
@@ III.2 Geodesic completion: massless particles
$aligned&f_{1}(x,y,z;E_{i})=\cos y\left(E_{1}\cos z-E_{2}\sin z\right)+E_{3}\sin y\,,\\ &f_{2}(x,y,z;E_{i})=\cos x\left(E_{1}\sin z+E_{2}\cos z\right)-\sin x\left[\sin y\left(E_{2}\sin z-E_{1}\cos z\right)+E_{3}\cos y\right],\\ &f_{3}(x,y,z;E_{i})=\cos x\left[\sin y\left(E_{2}\sin z-E_{1}\cos z\right)+E_{3}\cos y\right]+\sin x\left(E_{1}\sin z+E_{2}\cos z\right).$ (41)
@@ III.3 Geodesic completion: conformally coupled and massive particles
$aligned\dot{x}&=\cos y\frac{F_{1}\cos z-F_{2}\sin z+F_{3}\tan y}{a_{1}^{2}(t)\left(C\phi+m\right)}\\ &+\cos y\frac{\sin x\tan y(\cos x\sec y(F_{1}\sin z+F_{2}\cos z)-\sin x(-F_{1}\tan y\cos z+F_{2}\tan y\sin z+F_{3}))}{a_{2}^{2}(t)\left(C\phi+m\right)}\\ &-\cos y\frac{\cos x\tan y(\cos x(-F_{1}\tan y\cos z+F_{2}\tan y\sin z+F_{3})+\sin x\sec y(F_{1}\sin z+F_{2}\cos z))}{a_{3}^{2}(t)\left(C\phi+m\right)}$ (56)
$aligned\dot{y}&=\frac{-2\sin x\cos x\sin y\left(a^{2}_{2}(t)-a^{2}_{3}(t)\right)(F_{1}\cos z-F_{2}\sin z)+F_{3}\sin(2x)\cos y\left(a_{2}(t)^{2}-a_{3}(t)^{2}\right)}{a_{2}(t)^{2}a_{3}(t)^{2}\left(C\phi+m\right)}\\ &+\frac{2(F_{1}\sin z+F_{2}\cos z)\left(a_{2}(t)^{2}\sin^{2}x+a_{3}(t)^{2}\cos^{2}x\right)}{a^{2}_{2}(t)a^{2}_{3}(t)\left(C\phi+m\right)}$ (57)
$aligned\dot{z}&=\frac{\cos x(\cos x(-F_{1}\tan y\cos z+F_{2}\tan y\sin z+F_{3})+\sin x\sec y(F_{1}\sin z+F_{2}\cos z))}{a^{2}_{3}(t)\left(C\phi+m\right)}\\ &+\frac{\sin x(\sin x(-F_{1}\tan y\cos z+F_{2}\tan y\sin z+F_{3})-\cos x\sec y(F_{1}\sin z+F_{2}\cos z))}{a^{2}_{2}(t)\left(C\phi+m\right)}$ (58)
@@
$aligned\Delta\lambda_{\rm cp}&=\int^{T}_{0}dt\ \left[1+\frac{G_{1}^{2}}{a_{1}^{2}(t)}+\frac{G_{2}^{2}}{a_{2}^{2}(t)}+\frac{G_{3}^{2}}{a_{3}^{2}(t)}\right]^{-1/2}\\ &=\int^{T}_{0}\ dt\ \frac{1}{(m+C\phi)^{2}}\left[1+\frac{\check{f}_{1}^{2}}{a_{1}^{2}(t)}+\frac{\check{f}_{2}^{2}}{a_{2}^{2}(t)}+\frac{\check{f}_{3}^{2}}{a_{3}^{2}(t)}\right]^{-1/2}\,.$ (61)
@@
$aligned\Delta\lambda_{\rm cp}&=\left(m+\frac{C}{\sqrt{32\pi G}}\right)^{-2}\int_{0}^{T}dt\ \left[1+\frac{\check{f}_{1}^{2}}{a_{1}^{2}(t)}+\frac{\check{f}_{2}^{2}}{a_{2}^{2}(t)}+\frac{\check{f}_{3}^{2}}{a_{3}^{2}(t)}\right]^{-1/2}\\ &\simeq\left(m+\frac{C}{\sqrt{32\pi G}}\right)^{-2}\int_{0}^{T}dt\ \left[\frac{\check{f}^{2}_{2}a_{1}^{2}(t)a_{3}^{2}(t)}{\left(a_{1}(t)a_{2}(t)a_{3}(t)\right)^{2}}\right]^{-\frac{1}{2}}\\ &=\frac{1}{|\check{f}_{2}|}\left(m+\frac{C}{\sqrt{32\pi G}}\right)^{-2}\int_{0}^{T}t^{p_{2}}dt\propto T^{p_{2}+1}$ (62)
@@
$aligned\Delta\lambda_{\rm cp}&=\int^{T}_{0}dt\left[\frac{1}{S(t)}+\left(\frac{{G_{1}^{*}}^{2}}{a_{1}^{2}(t)}+\frac{{G_{2}^{*}}^{2}}{a_{2}^{2}(t)}+\frac{{G_{3}^{*}}^{2}}{a_{3}^{2}(t)}\right)\right]^{-1/2}\\ &=\int_{0}^{T}dt\left[\frac{1}{S(t)}+\frac{1}{\left(m+\frac{C}{\sqrt{32\pi GS(t)}}\right)^{2}}\frac{1}{S^{2}(t)}\left(\frac{{\check{f}_{1}}^{2}}{a_{1}^{2}(t)}+\frac{{\check{f}_{2}}^{2}}{a_{2}^{2}(t)}+\frac{{\check{f}_{3}}^{2}}{a_{3}^{2}(t)}\right)\right]^{-1/2}\,.$ (65)
@@
$aligned\Delta\lambda_{\rm cp}&\simeq\int_{0}^{T}dt\left[\left(1+\frac{L_{1}^{2n}}{t^{2n}}\right)^{-1}+\left(1+\frac{L_{1}^{2n}}{t^{2n}}\right)^{-2}\sum_{i=1}^{3}\check{f}_{i}^{2}a_{i}^{-2}(t)\right]^{-1/2}\\ &\simeq\int_{0}^{T}dt\left[\left(1+\frac{L_{1}^{2n}}{t^{2n}}\right)^{-1}+\left(1+\frac{L_{1}^{2n}}{t^{2n}}\right)^{-2}\sum_{i=1}^{3}\check{f}_{i}^{2}t^{-2p_{i}}\right]^{-1/2}\\ &\simeq\int_{0}^{T}dt\left[\left(1+\frac{L_{1}^{2n}}{t^{2n}}\right)^{-1}+\left(1+\frac{L_{1}^{2n}}{t^{2n}}\right)^{-2}\check{f}_{2}^{2}t^{-2p_{2}}\right]^{-1/2}$ (66)
@@ IV Taming chaos
[…]